# Control Pneumatic Soft Bending Actuator with Feedforward Hysteresis Compensation by Pneumatic Physical Reservoir Computing

### v1

@@ V Conclusion
This study introduced a novel PRC-based feedforward hysteresis compensation model, named the FPRC model, for controlling a pneumatic soft actuator. This proposed method used a dual-PAM pneumatic bending actuator as a physical reservoir, with a T-S fuzzy model processing the physical reservoir’s readout and generating the feedforward control command. The proposed FPRC model demonstrated comparable training outcomes to an ESN model while performing better in the test phase. The FPRC model also required less training time than the ESN model and showed a 97.6% reduction in the execution time cost during the test phase compared to the ESN model. Experiments validated the FPRC model’s effectiveness in controlling the soft actuator’s motion with both open-loop and closed-loop control system setups. The FPRC model’s strong robustness against real-world disturbances has also been experimentally verified.
The control systems in this work followed the model-free design, which is easily applicable yet effective. However, the absence of the soft actuator’s dynamic model makes investigations like the system stability challenging. Future work may consider using PRC to replace the ESN in online learning control systems [9, 19], where stability analyses have been discussed. Since removing the necessary components required by the PRC system, e.g., actuation and feedback units, may compromise its functionality, future works may focus on alleviating the PRC’s space requirements by miniaturizing the physical system or integrating the actuation and feedback units into a highly integrated architecture.

### v2

@@ V Conclusion
This study introduced a PRC-based feedforward hysteresis compensation method, termed the FPRC model, to control a pneumatic soft actuator’s bending motion. The proposed method used a dual-PAM pneumatic soft bending actuator as a physical reservoir, with a T-S fuzzy model processing the physical reservoir’s readout and generating the feedforward control command. The proposed FPRC model demonstrated comparable model training performance to an ESN model while performing better in the model test phase. The FPRC model required less training time than the ESN model and showed a 97.6% reduction in the execution time cost during the model test phase than the ESN. Experiments validated the FPRC model’s effectiveness in controlling the soft actuator’s motion with both open-loop and closed-loop control system applications. The FPRC model’s strong robustness against real-world disturbances has also been experimentally verified.
The control systems in this work followed the model-free design, which is easily applicable yet effective. However, the RC’s inherent complexity [20, 21] and the absence of the controlled soft actuator’s dynamic model make investigations like the system stability challenging. Future work could explore the use of PRC as a substitute for the ESN in online learning control systems [9, 21], where the stability analysis has been discussed. Additionally, as removing the necessary components required by the PRC system (e.g., actuation and feedback units) may compromise its functionality, future work could focus on reducing the system’s spatial footprint through miniaturization or integrating actuation and feedback units into a compact architecture. While replacing the pneumatic physical reservoir with more compact and scalable alternatives such as memristor-based PRCs [22] can help alleviate the system’s bulkiness, cared attention may be required in matching the PRC’s timescale with that of the controlled system–e.g., using a pneumatic soft body PRC for controlling a pneumatic soft actuator–to ensure the PRC’s computational performance [20].